Hello, my name is Lucy. I need help on a math problem. I have sort of eliminated the choices C and D but am still struggling on A and B. Do you think you can help? The math problem reads:
A company produced 10,288 crayons. They put 75 crayons in each large box they ship out. Which statement is true?
A. The company can fill 137 large boxes of crayons, with 11 crayons left over.
B. The company can fill 137 large boxes of crayons, with 13 crayons left over.
C. The company can fill 138 large boxes of crayons, with 10 crayons left over.
D. The company can fill 138 large boxes of crayons, with 14 crayons left over.

Answer by Alan3354(69443)

A company produced 10,288 crayons. They put 75 crayons in each large box they ship out. Which statement is true?
A. The company can fill 137 large boxes of crayons, with 11 crayons left over.
B. The company can fill 137 large boxes of crayons, with 13 crayons left over.
C. The company can fill 138 large boxes of crayons, with 10 crayons left over.
D. The company can fill 138 large boxes of crayons, with 14 crayons left over.
====================
10288/75 =~ 137.17333 full boxes
137*75 = 10275 in the boxes
-----
10288 - 10275 = 13 left over